# Event2Vec: Processing neuromorphic events directly by representations in vector space

### v1

@@ 5 Discussion
The neuromorphic event cameras bring new opportunities and challenges for computer vision. Researchers have made massive efforts to coordinate the gap between event-based vision and deep learning methods. In this article, we propose the first-of-its-kind event2vec representation and enable neural networks to process asynchronous events directly without losing any temporal resolution.
With the proposed event2vec representation, a vanilla transformer encoder with only 2 layers achieves impressive performance on the ASL-DVS classification task. The extremely small model size and high inference speed are more attractive than the test accuracy results. The parameter efficiency shows a promising prospect of deploying the model with event2vec on resource-limited edge computing devices. Although neuromorphic event cameras have $\mu s$-level temporal sensitivity, the computational speed of subsequent models cannot keep up with the preceding sensors in most cases. With the event2vec representation, the model gets rid of the latency of pre-processing the event data and has the potential to achieve $ms$-level response speed in an end-to-end scenario.
Beyond parameter efficiency, task accuracy, and running speed, the most promising feature of event2vec is that it aligns events to the domain of natural language processing, which has achieved significant progress and attracted interest even outside academia and industry. By combining event2vec representation with powerful large language models, various applications can be explored. For example, we can use a decoder-only transformer structure and train on events with self-supervised learning, then the model is expected to generate future events from previous events like an ”event-GPT”. We can also build multi-modal large models incorporating events, images, and audio for low-latency and robust autonomous driving under complex environments. In conclusion, our research has the potential to extend the research topics of neuromorphic computing to generative tasks and applications.

### v2

@@ 5 Discussion
The neuromorphic event cameras bring new opportunities and challenges for computer vision. Researchers have made massive efforts to bridge the gap between event-based vision and deep learning methods. In this article, we propose the first-of-its-kind event2vec representation and enable neural networks to process asynchronous events directly without losing any temporal resolution.
The results on classifying DVS Gesture and ASL-DVS datasets validate the feasibility of processing events directly, and preprocessing methods such as event-to-frame can be avoided. At the same time, with only a few randomly sampled events, the model achieves comparable accuracy to previous methods, which take all events as inputs. Fewer events lead to less computational costs, memory consumption, latency, and energy demands, showing a promising prospect of deploying the model with event2vec on resource-limited edge computing devices. Although neuromorphic event cameras have $\mu s$-level temporal sensitivity, the computational speed of subsequent models cannot keep up with the preceding sensors in most cases. With the event2vec representation, the model gets rid of the latency of pre-processing the event data and has the potential to achieve $ms$-level response speed in an end-to-end scenario.
The most promising feature of event2vec is that it aligns events to the domain of natural language processing, which has achieved significant progress and attracted interest even outside academia and industry. By combining event2vec representation with powerful large language models, various applications can be explored. For example, we can use a decoder-only transformer structure and train on events with self-supervised learning, then the model is expected to generate future events from previous events like an ”event-GPT”. We can also build multi-modal large models incorporating events, images, and audio for low-latency and robust autonomous driving under complex environments. In conclusion, our research has the potential to extend the research topics of neuromorphic computing to generative tasks and applications.